OpenStudy (anonymous):

X^2 = Y + a Y^2 = Z + a Z^2 = X + a - need help ??

OpenStudy (anonymous):

I need to find X,Y,Z

OpenStudy (anonymous):

By substitution: y = x^2-a <<<<this what I got (x^2-a)^2 = z+a so z = (x^2-a)^2-a then using z^2 = x+a we get ((x^2-a)^2-a)^2 = x+a Then working this out x^8-4x^6a+2x^4(3a^2-a)-4x^2(a^3-a^2)-x+a^4-2a^3+a^2-a=0 Also: x^2- x - a = 0 because x^2 = a + y so x^2 = a + sqrt(z + a) which exands to x^2 = a + sqrt(a + sqrt(a + x))

OpenStudy (anonymous):

Can someone check it if its right ??